OpenStudy (anonymous):

A bottling machine filled 3 of every 8 bottles with grape juice. The bottling machine filled 144 bottles altogether. How many of these bottles were filled with grape juice?

OpenStudy (anonymous):

3/8 of 144, or \[\frac{ 3 }{ 8}*144\]

OpenStudy (anonymous):

I have no idea how to multiply fractions. XD

OpenStudy (anonymous):

Put the 144 over 1\[\frac{ 3 }{ 8 }*\frac{ 144 }{ 1 }\] Then multiply straight across \[\frac{ 3*144 }{ 8*1}\]

OpenStudy (anonymous):

So...432 divided by 8?

OpenStudy (anonymous):

That's 54...

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you're welcome